12 Be sure to answer all parts. The decomposition of
Answer #1

The above reaction of decomposition of 2N2O5(g) is first order reaction, the Half life of first order reaction is as follows t1/2 = 0.693/K, in this reaction the rate constant is 0.00045s-1

t1/2 = 0.693/K

t1/2 = 0.693/0.00045

t1/2 = 1540

t1/2 = 1.54 X 10-3s